Lemon Ltd. offers executive training seminars using, in part, recorded lectures of a well-known speaker. The agreement calls for Lemon to pay a royalty for the use of the lectures. The lecturer's agent offers Lemon two options. The first option is revenue-based and Lemon agrees to pay 25 percent of its revenues to the speaker. The second option is a flat rate of $390,000 annually for the use of the lectures in these seminars. The royalty agreement will run one year and the royalty option chosen cannot be changed during the agreement. All other royalty terms are the same. Lemon charges $1,600 for the seminar and the variable costs for the seminar (excluding any royalty) is $300. Annual fixed costs (excluding any royalties) are $585,000. Required: a. What is the annual break-even level assuming: 1. The revenue-based royalty agreement? 2. The flat-rate royalty agreement? b. At what annual volume would the operating profit be the same regardless of the royalty option chosen? c. Assume an annual volume of 1,500 seminars. What is the operating leverage assuming: 1. The revenue-based royalty agreement? 2. The flat-rate royalty agreement? d. Assume an annual volume of 1,500 seminars. What is the margin of safety assuming: 1. The revenue-based royalty agreement? 2. The flat-rate royalty agreement?
Answer #1

Answer Page No 0 Revenue-based flat rate Selling price $1600 $1600 Variable cost ($300) ($300) 1 Royalty ($400) : (60) ContuiPage. No ① level = 650 seminars Revenue based, breals even ☺ Hat orate royally agreement, Break even level. $975000 $1300 = 1. $1350,000 1 Page No Operating income = (1500 seminars x $900) – 585000 . = 1350,000 - 585000 = $ 465000 Operating leverageTage. No 6 1. Operating deverage = al a manger og sately assuming devenue based a @ flat rate o Revenue besed vroyally agreem